# Birds look like cars: Adversarial analysis of intrinsically interpretable deep learning

### v1

@@ Abstract
A common belief is that intrinsically interpretable deep learning models ensure a correct, intuitive understanding of their behavior and offer greater robustness against accidental errors or intentional manipulation. However, these beliefs have not been comprehensively verified, and growing evidence casts doubt on them. In this paper, we highlight the risks related to overreliance and susceptibility to adversarial manipulation of these so-called “intrinsically (aka inherently) interpretable” models by design. We introduce two strategies for adversarial analysis with prototype manipulation and backdoor attacks against prototype-based networks, and discuss how concept bottleneck models defend against these attacks. Fooling the model’s reasoning by exploiting its use of latent prototypes manifests the inherent uninterpretability of deep neural networks, leading to a false sense of security reinforced by a visual confirmation bias. The reported limitations of prototype-based networks put their trustworthiness and applicability into question, motivating further work on the robustness and alignment of (deep) interpretable models.
Figure 1: Birds look like cars. Interpretation of image recognition by a ProtoViT classifying bird species based on car prototypes with 85% accuracy. We overemphasize that what looks alike to humans is not representative of what is similar according to the model. Manipulating the model’s reasoning by exploiting its use of latent prototypes manifests the inherent uninterpretability of prototype-based networks, which may be masked by human overreliance due to visual confirmation bias.
@@ 1 Introduction
Deep learning has seen remarkable development over the last decade, solving with superhuman efficiency a number of important tasks, such as object classification, segmentation, detection, etc. Unfortunately, the increasing complexity of (deep) machine learning models poses a challenge to their reliable adoption in high-stakes decisions like in medicine (Rudin, 2019; Combi et al., 2022; Jia et al., 2022; Baniecki et al., 2025b) or security (Andresini et al., 2022; Al-Essa et al., 2024). Even beyond automated decision-making, deep learning models allow acquiring insightful knowledge that leads to valuable discoveries, e.g. discovery of a structural class of antibiotics with explainable deep learning (Wong et al., 2024), better understanding of optimal strategies in board games (McGrath et al., 2022) or formulation of the laws of physics (Schmidt and Lipson, 2009). A limitation to such future discoveries is the black-box nature of deep learning architectures. To this end, a wide spectrum of interpretable machine learning methods were proposed, ranging from advances in intrinsically (aka inherently) interpretable models (Chen et al., 2008; Caruana et al., 2015; Arik and Pfister, 2021; Zhong et al., 2023) to improvements in post-hoc explainability algorithms for understanding large and complex models (Biecek and Burzykowski, 2021; McGrath et al., 2022; Covert et al., 2024; Taimeskhanov et al., 2024; Baniecki et al., 2025a; Chrabaszcz et al., 2025).
@@
But post-hoc explanations only approximate the behavior of the model and can be inaccurate (Rudin, 2019; Covert et al., 2024; Hesse et al., 2024). To ensure that the model and explanations are consistent, various interpretable deep learning architectures are proposed (Li et al., 2022) like prototype-based networks (Chen et al., 2019), bag-of-local-features (Brendel and Bethge, 2019), concept bottleneck models (Koh et al., 2020), or most recently B-cos networks (Böhle et al., 2024), which claim to offer image classification accuracy on par with their uninterpretable counterparts. The growing popularity of these model families comes from the intuitive, though never comprehensively tested, assumptions that inherently interpretable models show explanations consistent with human intuition, and are robust to accidental errors or intentional manipulation. After all, could these assumptions be wrong?
@@
Hoffmann et al. (2021) was the first to question the intrinsic interpretability of prototype-based networks relying on similarities between embeddings of uninterpretable network backbones. Ramaswamy et al. (2023) highlight that the choice of the probe distribution for concept-based explainability has a profound impact on the model’s interpretation. Based on their objections, we designed a set of examples that put into question the interpretability of prototypes. For example, Figure 1 shows an interpretation of the prototype-based network (specifically ProtoViT, Ma et al., 2024), where an image of a bird is classified based on features extracted from car images. Our proposed prototype manipulation highlights that visual confirmation bias (Klayman, 1995; Kim et al., 2022) is a threat potentially masking these models’ inherent uninterpretability. Furthermore, the manipulation might be hard to discover when only small patches and prototypes are presented.
@@
Aside from the problem with interpretability, we also still know little about the adversarial robustness of prototype-based networks. While there is a growing number of evidence that deep models are fragile, from the first adversarial examples fooling neural networks (Szegedy et al., 2014), through extensive robustness benchmarks for image classification (Hendrycks and Dietterich, 2019), also specifically in medical applications (Finlayson et al., 2019; Jia et al., 2022), to manipulation of large foundation models (Jankowski et al., 2024). Relatively well studied are adversarial threats to post-hoc explainability like adversarial examples and backdoor attacks against saliency maps (Noppel and Wressnegger, 2024), or data poisoning attacks against feature importance and effects (Baniecki and Biecek, 2024). A natural question arises: Are intrinsically interpretable deep learning models robust and secure?
@@
Our adversarial analysis aims to motivate further improvements in the alignment and security of deep interpretable models. Although adversarial examples crafted with projected gradient descent were already used to evaluate prototype stability (Huang et al., 2023a) and spatial misalignment (Sacha et al., 2024), as well as the robustness of concept bottleneck models (Sinha et al., 2023), we consider a threat of backdoor attacks instead (Noppel et al., 2023). To illustrate this, Figure 2 shows an adversarially manipulated interpretation of the prototype-based network (specifically PIP-Net, Nauta et al., 2023b), where an image of a skin lesion is classified as either benign or malignant. We found it surprising that it is possible to embed a backdoor in a deep interpretable model in a way to conceal the adversarial trigger, or even manipulate the interpretation to contradict the prediction.
@@
Figure 2: Backdoor attacks on a PIP-Net model classifying malignant skin lesions with 85% accuracy. The naive backdoor attack could be detected by a visual inspection of explanations where one prototype highlights the trigger. We exploit the vulnerabilities of latent-based models considering two adversarial scenarios: disguising the attack when the model provides an original explanation for a new prediction, and a red herring that further manipulates the explanation, naturally covering up the reason why the prediction changed from benign to malignant.
@@
We hope that this critical analysis will identify, but also help to improve, weak spots in the design of deep interpretable models. Supplementary code for methods and experiments will be available at https://github.com/hbaniecki/adversarial-analysis-of-interpretable-dl.
2 Related work
2.1 Interpretable deep learning models
@@
Prototype-based networks are applied to medical image classification with the motivation to support alignment and correction of deep learning models, e.g. in histopathology (Rymarczyk et al., 2022a), dermatology and radiology (Nauta et al., 2023a), or even detection of Alzheimer in 3D magnetic resonance images (Santi et al., 2024b). Prototype-based networks have been adapted to text classification (Fanconi et al., 2023; Hong et al., 2023). Beyond vision and language, PivotTree (Cascione et al., 2024) proposes using pivots as whole inputs in a data-agnostic manner instead of prototypical parts, allowing to experiment with different data modalities like tables and time series. Most recently, in this line of research, the case-based paradigm is adapted for vision transformers in ProtoPFormer (Xue et al., 2024) and ProtoViT (Ma et al., 2024) architectures.
@@
2.2 Adversarial analysis of interpretability & explainability
Ante-hoc interpretability.
The general consensus is that inherent (ante-hoc) interpretability entails a higher degree of safety in machine learning (Otte, 2013; Rudin, 2019; Wei et al., 2022). To the best of our knowledge, Hoffmann et al. (2021) were the first to question this assumption, showing how a simple adversarial example and data poisoning attack can break ProtoPNet. Adversarial examples prove useful to evaluate the stability and (mis)alignment of prototype-based networks (Huang et al., 2023a; Sacha et al., 2024). Sourati et al. (2024) stress-test prototype-based networks for interpretable text classification using several text-specific perturbation attacks, as well as show that adversarial augmented training improves the robustness of ProtoPNets. Similarly, Sinha et al. (2023) examine the robustness of CBMs to input perturbations and propose adversarial training to defend against such a threat. We provide further evidence countering the above-mentioned belief by focusing on state-of-the-art architectures, prototype manipulation, and safety-critical backdoor attacks.
Our research relates to the evaluation of interpretable deep learning models beyond the notions of robustness and security. Kim et al. (2022) perform a large-scale human evaluation of visual explanations comparing prototype-based networks to the popular post-hoc approach Grad-CAM (Taimeskhanov et al., 2024). It provides concrete insights into the issue of visual confirmation bias in interpretability. Xu-Darme et al. (2023) evaluate the faithfulness and relevance of different methods for generating patch visualization from similarity maps in ProtoPNet and ProtoTree. Bontempelli et al. (2023) propose ProtoPDebug, a framework for correcting and aligning prototype-based networks with human-feedback. Ramaswamy et al. (2023) highlight that the choice of the probe dataset has a profound impact on the generated concept-based explanations, including CBMs. We obtain similar results for prototype-based networks.
Post-hoc explainability.
3 Methods
3.1 Preliminaries
@@
The final classification is based on the highest output product $\mathbf{p}\cdot\theta_{h}$ per class. During training, additional sparsity regularization is added with $\log\left(1+\left(\mathbf{p}\cdot\theta_{h}\right)^{2}\right)$. The model is trained end-to-end with classification loss $\mathcal{L}_{\mathrm{C}}$, e.g. negative log-likelihood loss. PIP-Net proposes a contrastive pre-training of prototypes (similar to Wang and Isola, 2020) with the combined loss function $\mathcal{L}(\theta)=\lambda_{\mathrm{C}}\mathcal{L}_{\mathrm{C}}+\lambda_{\mathrm{A}}\mathcal{L}_{\mathrm{A}}+\lambda_{\mathrm{U}}\mathcal{L}_{\mathrm{U}}$. Alignment loss $\mathcal{L}_{\mathrm{A}}(\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime})=-\sum_{i}\log(\mathbf{z}^{\prime}_{i}\cdot\mathbf{z}^{\prime\prime}_{i})$ computes the similarity between latent representations of two augmented views of input $\mathbf{x}$ as their dot product. Uniformity loss $\mathcal{L}_{\mathrm{U}}(\mathbf{p})=-\sum_{i}\log\left(\tanh\left(\sum_{\mathbf{x}\in\text{batch}}\mathbf{p}_{i}^{\mathbf{x}}\right)\right)$ optimizes to learn non-trivial representations that make use of the whole prototype space. For input $\mathbf{x}$, parts of its latent representations $\mathbf{z}$ corresponding to the closest prototypes $\mathbf{p}$, or most contributing ones $\mathbf{p}\cdot\theta_{h}$, can be visualized as heatmaps overlayed over the input (Fig. 2). Furthermore, the heatmaps can be converted to bounding boxes denoting important input patches (Fig. 7), where one can interpret each prototype by visualizing parts of other inputs (e.g. from $\mathcal{D}_{\mathrm{train}}$).
@@
where $\theta_{g,i}$ and $\mathbf{z}_{j}$ correspond to the latent representation of up to $t$ tokens (patches). Thus, Figure 1 shows similarity scores between $-4$ (practically 00) and $4$ for $t=4$. Similar to the original ProtoPNet (Chen et al., 2019), training ProtoViT includes an important prototype projection stage where each stored embedding $\theta_{g,j}$ is substituted with the latent representation $\mathbf{z}$ of the closest input parts from $\mathcal{D}_{\mathrm{train}}$ as measured by the summed cosine similarity (akin Eq. 1). The so-called prototypes $\theta_{g}$ are then frozen, and the last layer $h$ is further fine-tuned and regularized to achieve sparse, interpretable prediction. Contrary to PIP-Net, the projection aims to ground the interpretable representation $\mathbf{p}$ in latent representations $\theta_{g}$ that can be visualized as “cases” from $\mathcal{D}_{\mathrm{train}}$ in the model’s reasoning process (see Fig. 6A).
3.2 Adversarial analysis of prototype-based networks
@@
In line with related work on adversarial examples (Ghorbani et al., 2019; Hoffmann et al., 2021; Sacha et al., 2024), model manipulation (Heo et al., 2019) and backdoor attacks (Noppel et al., 2023), we consider a white-box attacker or robustness evaluator with full access to the data and model for critically analyzing the interpretability of state-of-the-art architectures and improving defenses. In prototype manipulation, one specifically targets $\theta_{g}$ and aims to fine-tune the already trained model. Executing a backdoor attack requires an adversary to influence the training process and/or data.
Prototype manipulation.

### v2

@@ Abstract
A common belief is that intrinsically interpretable deep learning models ensure a correct, intuitive understanding of their behavior and offer greater robustness against accidental errors or intentional manipulation. However, these beliefs have not been comprehensively verified, and growing evidence casts doubt on them. In this paper, we highlight the risks related to overreliance and susceptibility to adversarial manipulation of these so-called “intrinsically (aka inherently) interpretable” models by design. We introduce two strategies for adversarial analysis with prototype manipulation and backdoor attacks against prototype-based networks, and discuss how concept bottleneck models defend against these attacks. Fooling the model’s reasoning by exploiting its use of latent prototypes manifests the inherent uninterpretability of deep neural networks, leading to a false sense of security reinforced by a visual confirmation bias. The reported limitations of part-prototype networks put their trustworthiness and applicability into question, motivating further work on the robustness and alignment of (deep) interpretable models.
Figure 1: Birds look like cars. Interpretation of image recognition by a ProtoViT classifying bird species based on car prototypes with 85% accuracy. We overemphasize that what looks alike to humans is not representative of what is similar according to the model. Below is a schema of the prototype-based model’s architecture with a highlighted point of failure. Manipulating the model’s reasoning by exploiting its use of latent prototypes ($\theta_{g}$) manifests the inherent uninterpretability of prototype-based networks, which may be masked by human overreliance due to visual confirmation bias.
@@ 1 Introduction
Deep learning has seen remarkable development over the last decade, solving with superhuman efficiency a number of important tasks, such as object classification, segmentation, detection, etc. Unfortunately, the increasing complexity of (deep) machine learning models poses a challenge to their reliable adoption in high-stakes decisions like in medicine (Rudin, 2019; Combi et al., 2022; Jia et al., 2022; Baniecki et al., 2025b) or security (Andresini et al., 2022; Al-Essa et al., 2024). Even beyond automated decision-making, deep learning models allow acquiring insightful knowledge that leads to valuable discoveries, e.g. discovery of a structural class of antibiotics with explainable deep learning (Wong et al., 2024), better understanding of optimal strategies in board games (McGrath et al., 2022) or formulation of the laws of physics (Schmidt and Lipson, 2009). A limitation to such future discoveries is the black-box nature of deep learning architectures. To this end, a wide spectrum of interpretable machine learning methods was proposed, ranging from advances in intrinsically (aka inherently) interpretable models (Chen et al., 2008; Caruana et al., 2015; Arik and Pfister, 2021; Zhong et al., 2023) to improvements in post-hoc explainability algorithms for understanding large and complex models (Biecek and Burzykowski, 2021; McGrath et al., 2022; Covert et al., 2024; Taimeskhanov et al., 2024; Baniecki et al., 2025a; Chrabaszcz et al., 2025).
@@
But post-hoc explanations only approximate the behavior of the model and can be inaccurate (Rudin, 2019; Covert et al., 2024; Hesse et al., 2024). To ensure that the model and explanations are consistent, various interpretable deep learning architectures are proposed (Li et al., 2022) like prototype-based networks (Chen et al., 2019), bag-of-local-features (Brendel and Bethge, 2019), concept bottleneck models (Koh et al., 2020), or most recently B-cos networks (Böhle et al., 2024), which claim to offer image classification accuracy on par with their uninterpretable counterparts. The growing popularity of these model families comes from the intuitive, though never comprehensively tested, assumptions that inherently interpretable models show explanations consistent with human intuition, and are robust to accidental errors or intentional manipulation.
After all, could these assumptions be wrong?
@@
Hoffmann et al. (2021) was the first to question the intrinsic interpretability of prototype-based networks relying on similarities between embeddings of uninterpretable network backbones. Ramaswamy et al. (2023) highlight that the choice of the probe distribution for concept-based explainability has a profound impact on the model’s interpretation. Based on their objections, we designed a set of examples that put into question the interpretability of prototypes. For example, Figure 1 shows an interpretation of the prototype-based network (specifically ProtoViT, Ma et al., 2024), where an image of a bird is classified based on features extracted from car images. Our proposed prototype manipulation highlights that visual confirmation bias (Klayman, 1995; Kim et al., 2022) is a threat, potentially masking these models’ inherent uninterpretability. Furthermore, the manipulation might be hard to discover when only small patches and prototypes are presented.
@@
Aside from the problem with interpretability, we also still know little about the adversarial robustness of prototype-based networks. While there is a growing body of evidence that deep models are fragile, from the first adversarial examples fooling neural networks (Szegedy et al., 2014), through extensive robustness benchmarks for image classification (Hendrycks and Dietterich, 2019), also specifically in medical applications (Finlayson et al., 2019; Jia et al., 2022), to manipulation of large foundation models (Jankowski et al., 2024). Relatively well-studied are adversarial threats to post-hoc explainability, like adversarial examples and backdoor attacks against saliency maps (Noppel and Wressnegger, 2024), or data poisoning attacks against feature importance and effects (Baniecki and Biecek, 2024).
A natural question arises: Are intrinsically interpretable deep learning models robust and secure?
@@
Our adversarial analysis aims to motivate further improvements in the alignment and security of deep interpretable models. Although adversarial examples crafted with projected gradient descent were already used to evaluate prototype stability (Huang et al., 2023a) and spatial misalignment (Sacha et al., 2024), as well as the robustness of concept bottleneck models (Sinha et al., 2023), we consider a threat of backdoor attacks instead (Noppel et al., 2023). To illustrate this, Figure 2 shows an adversarially manipulated interpretation of the prototype-based network (specifically PIP-Net, Nauta et al., 2023b), where an image of a skin lesion is classified as either benign or malignant. We found it surprising that it is possible to embed a backdoor in a deep interpretable model in a way that conceals the adversarial trigger, or even manipulates the interpretation to contradict the prediction.
@@
Figure 2: Backdoor attacks on a PIP-Net model classifying malignant skin lesions with 85% accuracy. The naive backdoor attack could be detected by a visual inspection of explanations, where one prototype highlights the trigger. Below is a schema of the prototype-based model’s architecture with a highlighted point of failure. We exploit the vulnerabilities of latent-based models ($\theta_{f}$), considering two adversarial scenarios: disguising the attack when the model provides an original explanation for a new prediction, and a red herring that further manipulates the explanation, naturally covering up the reason why the prediction changed from benign to malignant.
@@
We hope that this critical analysis will identify, but also help to improve, failure points in the design of deep interpretable models. The structure of this paper is as follows. We first provide the necessary background on intrinsically interpretable deep learning with an introduction to our notation in Section 2. Section 3 discusses work related to adversarial analysis of ante-hoc interpretability and post-hoc explainability. We introduce prototype manipulation in Section 4.1 and backdoor attacks in Section 4.2; then, we evaluate them in Sections 5.1 and 5.2, respectively. Finally, we summarize the insights from our empirical analysis and also related studies in Section 6. Supplementary code for methods and experiments is available at https://github.com/hbaniecki/adversarial-analysis-of-interpretable-dl.
2 Background and notation
@@
Prototype-based networks are applied to medical image classification with the motivation to support alignment and correction of deep learning models, e.g. in histopathology (Rymarczyk et al., 2022a), dermatology and radiology (Nauta et al., 2023a), or even detection of Alzheimer’s in 3D magnetic resonance images (Santi et al., 2024b). Prototype-based networks have been adapted to text classification (Fanconi et al., 2023; Hong et al., 2023). Beyond vision and language, PivotTree (Cascione et al., 2024) proposes using pivots as whole inputs in a data-agnostic manner instead of prototypical parts, allowing for experimentation with different data modalities like tables and time series. Most recently, in this line of research, the case-based paradigm is adapted for vision transformers in ProtoPFormer (Xue et al., 2024) and ProtoViT (Ma et al., 2024) architectures.
@@
The final classification is based on the highest output product $\mathbf{p}\cdot\theta_{h}$ per class, where during training, additional sparsity regularization is added with $\log\left(1+\left(\mathbf{p}\cdot\theta_{h}\right)^{2}\right)$. The model is trained end-to-end with classification loss $\mathcal{L}_{\mathrm{C}}$, e.g. negative log-likelihood loss. PIP-Net proposes a contrastive pre-training of prototypes (similar to Wang and Isola, 2020) with the combined loss function
$\mathcal{L}(\theta)=\lambda_{\mathrm{C}}\mathcal{L}_{\mathrm{C}}+\lambda_{\mathrm{A}}\mathcal{L}_{\mathrm{A}}+\lambda_{\mathrm{U}}\mathcal{L}_{\mathrm{U}},$ (1)
where alignment loss
$\mathcal{L}_{\mathrm{A}}(\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime})=-\sum_{i}\log(\mathbf{z}^{\prime}_{i}\cdot\mathbf{z}^{\prime\prime}_{i})$ (2)
computes the similarity between latent representations of two augmented views of input $\mathbf{x}$ as their dot product. In what follows, uniformity loss
$\mathcal{L}_{\mathrm{U}}(\mathbf{p})=-\sum_{i}\log\left(\tanh\left(\sum_{\mathbf{x}\in\text{batch}}\mathbf{p}_{i}^{\mathbf{x}}\right)\right)$ (3)
optimizes to learn non-trivial representations that make use of the whole prototype space. For input $\mathbf{x}$, parts of its latent representations $\mathbf{z}$ corresponding to the closest prototypes $\mathbf{p}$, or most contributing ones $\mathbf{p}\cdot\theta_{h}$, can be visualized as heatmaps overlayed over the input (Fig. 2). Furthermore, the heatmaps can be converted to bounding boxes denoting important input patches (Fig. 7), where one can interpret each prototype by visualizing parts of other inputs, e.g. from $\mathcal{D}_{\mathrm{train}}$.
@@
where $\theta_{g,i}$ and $\mathbf{z}_{j}$ correspond to the latent representation of up to $t$ tokens (patches). Thus, Figure 1 shows similarity scores between 4-4 (practically 0) and $4$ for $t=4$. Similar to the original ProtoPNet (Chen et al., 2019), training ProtoViT includes an important prototype projection stage where each stored embedding $\theta_{g,j}$ is substituted with the latent representation $\mathbf{z}$ of the closest input parts from $\mathcal{D}_{\mathrm{train}}$ as measured by the summed cosine similarity (akin Eq. 4). The so-called prototypes $\theta_{g}$ are then frozen, and the last layer $h$ is further fine-tuned and regularized to achieve sparse, interpretable predictions. Prototype projection is a single-stage process (cf. Ma et al., 2024, figure 5). Contrary to PIP-Net, the projection aims to ground the interpretable representation $\mathbf{p}$ in latent representations $\theta_{g}$ that can be visualized as “cases” from $\mathcal{D}_{\mathrm{train}}$ in the model’s reasoning process (see Fig. 6A).
3 Related work
Ante-hoc interpretability.
The general belief is that inherent (ante-hoc) interpretability entails a higher degree of safety in machine learning (Otte, 2013; Rudin, 2019; Wei et al., 2022). To the best of our knowledge, Hoffmann et al. (2021) were the first to question this assumption, showing how a simple adversarial example and data poisoning attack can break ProtoPNet. Adversarial examples prove useful to evaluate the stability and (mis)alignment of prototype-based networks (Huang et al., 2023a; Sacha et al., 2024). Sourati et al. (2024) stress-test prototype-based networks for interpretable text classification using several text-specific perturbation attacks, as well as show that adversarial augmented training improves the robustness of ProtoPNets. Similarly, Sinha et al. (2023) examine the robustness of CBMs to input perturbations and propose adversarial training to defend against such a threat. Contrary to Hoffmann et al. (2021) and Sinha et al. (2023), we analyse the ProtoViT and PIP-Net architectures, proposing alternative manipulation techniques. We provide further evidence countering the above-mentioned belief by focusing on state-of-the-art architectures, prototype manipulation, and safety-critical backdoor attacks.
Our research relates to the evaluation of interpretable deep learning models beyond the notions of robustness and security. Kim et al. (2022) perform a large-scale human evaluation of visual explanations comparing prototype-based networks to the popular post-hoc approach Grad-CAM (Taimeskhanov et al., 2024). It provides concrete insights into the issue of visual confirmation bias in interpretability. Xu-Darme et al. (2023) evaluate the faithfulness and relevance of different methods for generating patch visualization from similarity maps in ProtoPNet and ProtoTree. Bontempelli et al. (2023) propose ProtoPDebug, a framework for correcting and aligning prototype-based networks with human-feedback. Such an interactive approach could potentially be leveraged as a countermeasure against our proposed and related manipulation techniques. Ramaswamy et al. (2023) highlight that the choice of the probe dataset has a profound impact on the generated concept-based explanations, including CBMs. Contrary, we show this impact in prototype-based networks, and go further to demonstrate the possibility of hiding explanation-aware backdoors in deep interpretable models.
Post-hoc explainability.
4 Methods
@@
In line with related work on adversarial examples (Ghorbani et al., 2019; Hoffmann et al., 2021; Sacha et al., 2024), model manipulation (Heo et al., 2019), and backdoor attacks (Noppel et al., 2023), we consider a white-box attacker or robustness evaluator with full access to the data and model for critically analyzing the interpretability of state-of-the-art architectures and improving defenses. In prototype manipulation, one specifically targets $\theta_{g}$ and aims to fine-tune the already trained model. Note that prototype images and their parts are an inherent element of the model, like its weights, accessible to any potential user at inference time. Executing a backdoor attack requires an adversary to influence the training process and/or data to modify $\theta_{f}$. Both methods can be considered training-time attacks, while using the backdoor additionally requires modifying an input to the model at inference time.
4.1 Prototype manipulation